A body starting from rest accelerates at a constant rate m/sfor some time after which it retards at a constant rate b m/s2 to come to rest finally. If the total time elapsed is t sec,the maximum velocity attended by a body is given by

Grade:11

1 Answers

Vikas TU
7 years ago
If u make the graph of velocity vs x in time intervals t1 and t2 respectively,
then u would get 
t1 + t2 = T
or
v/a + v/b = T
v = (ab)*T/(a+b)